A soft drink machine outputs a mean of 28ounces per cup. The machine's output is normally distributed with a standard deviation of 3 ounces. What is the probability of putting less than 30 ounces in a cup? Round your answer to four decimal places.
Answer #1

Here it is given that distribution is normal with mean=28 and standard deviation=3

We need to find P (x<30)

As distribution is normal we can convert x to z

P(x<30)=P(z<\frac{30-28}{3}) = P(z<0.67)=0.7486